OpenStudy (automaticloveletter):

How many hours will Sherwin most likely sleep if he consumes 9 cups of coffee?

OpenStudy (automaticloveletter):

Link to graph: http://prntscr.com/79a8hp Answers: 1, because y = −x + 10 2, because y = −x + 10 9, because y = −x + 10 10, because y = −x + 10 I chose A (1, because y = -x + 10) Am I right or wrong?

OpenStudy (perl):

that is correct